OpenStudy (anonymous):

Rennie kicks a field goal with an initial vertical velocity of 31 meters per second. How long will it take the football to hit the ground?

OpenStudy (anonymous):

um sorry i dont no this one but does it give you any minutes or something?

OpenStudy (anonymous):

–4.9t2 + 31t = 0 –4.9t2 – 31t = 0 16t2 + 31t = 0 –16t2+ 31t = 0

OpenStudy (anonymous):

one of those are the answer

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

okay, thanks!

OpenStudy (anonymous):

A